OpenStudy (anonymous):

Alright this is a continuation of my regular tetrahedron inside a cube problem, after finding the measure of the angle between the carbon molecules and the hydrogen molecules in methane. Use dot product to find the angle between two adjacent edges (sharing a common vertex) of the tetrahedron; and the angle between two opposite edges. Explain your answer using symmetry.

OpenStudy (anonymous):

here is our tetrahedron \[(1, 1, 1), (-1, 1, -1), (-1, -1, 1), (1, -1, -1)\]

OpenStudy (anonymous):

actually, those vectors represent the vertices of that tetrahedron. What do I do if I want to find vectors that represent the edges of my tetrahedron ?

OpenStudy (anonymous):

vector addition. right?

OpenStudy (anonymous):

what is the vector from (1,1,1) to (-1,1,-1)?

OpenStudy (anonymous):

their difference, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[A := (1,1,1)\] \[B:=(-1,1,-1)\] \[C:=(-1,-1,1)\] \[D:=(1,-1,-1)\]

OpenStudy (anonymous):

\[AB = A - B = (1,1,1)-(-1,1,-1)=(2,0,2)\]

OpenStudy (anonymous):

umm or is it the other way around? B-A?

OpenStudy (anonymous):

(-1,1,-1)-(1,1,1)=(-2,0,-2)=AB

OpenStudy (anonymous):

yeah it's the other way around \[AB = B-A = (-2,0,2)\]

OpenStudy (anonymous):

\[AC = C-A=(-1,-1,1)-(1,1,1)=(-2,-2,0)\] \[AD = D-A=(1,-1,-1)-(1,1,1)=(0,-2,-2)\]

OpenStudy (anonymous):

those are like 3 adjacent edges and I only need 2 to solve the problem.

OpenStudy (anonymous):

What about if I want to find the angle of 2 *opposite* edges of a tetrahedron? what are opposite edges?

OpenStudy (anonymous):

btw AB should be \[(-2,0,-2)\]

OpenStudy (anonymous):

hmm when you visualize it.... the angle between 2 opposite edges of a tetrahedron is pi/2 radians.

OpenStudy (anonymous):

1 more problem

OpenStudy (anonymous):